OpenStudy (anonymous):

2y+x=4 3y-2x=13 x=-2,y=3 Just trying to plot these lines on a linear graph. I come up with the right x and y values for the second equasion- y=2x/3+13/3 but when I when I make a table of values for the first equation, I keep coming up with x=-2,y=1 based on y=x/2+2. Can anyone help?

OpenStudy (anonymous):

y=-1/2 x +2 y=2/3x +13/3

OpenStudy (anonymous):

So plot these two lines

OpenStudy (anonymous):

cheers... i had that on the first attempt but must of done something stupid.....again! hang on, i didn't have the negative gradient. I knew the gradient was a 1/2 and the y intercept was 2...

OpenStudy (anonymous):

Yeah, when you moved the x to other side, it should have been -x

OpenStudy (anonymous):

2y+x=4 multiply this by 2 we get 4y+2x=8 3y-2x=13 3y-2x=13 add these two together ----------- 7y = 21 y=3 ans........................subst this to eq 1 we get 2(3)+x=4 x= 4-6=-2 ans.. so the ans is (-2,3)=(x,y)

OpenStudy (anonymous):

id21....that wolfram program looks amazing!

OpenStudy (id21):

You can use it by yourself! :) It is online http://www.wolframalpha.com

OpenStudy (anonymous):

To plot a bunch of function: example 2x and x^2 from 0 to 2 Plot[{2x,x^2},{x,0,2}]

OpenStudy (id21):

I used this syntax: plot y=2x/3+13/3 and y=-x/2+2 from -4 to 4